# Towards grid-aware multi-period flexibility aggregation – A constrained zonotope approach

### v1

@@ V Numerical Case Study
All numerical results in this section are performed on a system with an eight core and 16 thread Ryzen 9800X3D CPU paired with 64GB of 6000MHz RAM. Provided computation times are the average of ten executions. We use JuMP.jl [23] and MOSEK [24] to solve optimization problems. For convex sets, Polyhedra.jl allows to extract the set directly form the JuMP.jl model as polytope in H-representation. For the problem formulation, we consider LinDistFlow with and without loss linearization as as discussed in Section II. This allows to achieve a convex approximation while maintaining the crucial information of reactive power and voltage magnitude in the DSO grid model. As the underlying problem, we consider a multi-stage Optimal Power Flow (MS-OPF) problem, which allows flexibililty aggregation across multiple time steps and serves to evaluate computational scalability for horizons of up to $N=96$.
@@ V-A Comparison of Computation Times
First, we consider a small-scale, 4-bus radial distribution grid adapted from MATPOWER, called case4dist [25]. The grid has been extended to include two additional renewable generators at buses three and four, thereby increasing the system’s overall active and reactive power flexibility. This small test case allows the comparison of the presented method with the baseline approach of polytope projection, which can still be evaluated over a short varying time horizon of the MS-OPF. While this projection-based approach does not scale well to larger grids, it remains suitable for evaluating performance in this compact setting.
TABLE I: Computation times of the presented method and polytope projection on the 4-bus distribution grid model for up to $N=4$ timesteps. Fastest computation times in green.
Horizon NN Approach Offline [s] Online [s] Total [s]
1 $\mathcal{CZ}$ 0.000 657 250.000\,657\,25 $1.429\,99\text{\times}{10}^{-6}$ 0.000 658 650.000\,658\,65
[…]
2 $\mathcal{CZ}$ 0.002 958 739 90.002\,958\,739\,9 $7.8000\text{\times}{10}^{-7}$ 0.002 959 520.002\,959\,52
[…]
3 $\mathcal{CZ}$ 0.007 280 230.007\,280\,23 $1.02\text{\times}{10}^{-6}$ 0.007 281 250.007\,281\,25
[…]
4 $\mathcal{CZ}$ 0.011 396 410 000 000 0010.011\,396\,410\,000\,000\,001 $29\,999\,999\,999\,999\,9\text{\times}{10}^{-6} 0.011$ 397 440.011\,397\,44
@@
V-B Evaluation on Medium-sized Feeder
[…]
@@ V-B Evaluation on Medium-sized Feeder
First, we demonstrate the equivalence of the reformulation of the feasible set as $\mathcal{CZ}$. Figure 4 shows that approximating the variable bounds does not impact the accuracy of the resulting $\mathcal{CZ}$. The FOR modeled as $\mathcal{CZ}$ $\mathcal{F}_{\mathcal{CZ}}^{LDF-LL}$ is equivalent to the respective FOR modeled as polytope $\mathcal{F}^{LDF-LL}$. As described in Subsection III-B, this holds as long as the bounding zonotope over-approximates the polytope and all constraints defining the polytope are added to the $\mathcal{CZ}$ representation in the next step of the conversion. Hence, all following computations use sufficiently large variable bounds to decrease computation time.
@@
Table II shows the respective timeseries of required computation times for the reformulation of the feasible set as $\mathcal{CZ}$. For the given results, the constraint intersection is performed in parallel. The results clearly show that despite the fact that the computation time scales noticeably, it remains suitable for offline computation on the used test problem, even for a considered time horizon length of $N=96$. Further, $\mathcal{CZ}$ projection can be performed in sub-second times on the 15-bus distribution grid case, even for large time horizons $N$. This highlight the suitability for the presented method to reduce complexity in ADP. Paired with acceptable conversion computation times, this allows the application of ADP to considerably larger problem sizes, by using the presented method to model the feasible sets of the sub problems as $\mathcal{CZ}$.
@@ VI Conclusion and Future Work
This paper addresses the central computational bottleneck arising in ADP, which is the efficient computation of set projections. By utilizing $\mathcal{CZ}\mathrm{s}$ as beneficial set representations we can significantly reduce the computation times required for flexibility aggregation via set projection of convex feasible sets. This allows to perform time-dependent flexibility aggregation in the setting of multi-stage OPF and on large-scale problems without loss of accuracy. Further we show that most of the computation can be performed offline and set modifications can be considered efficiently.
Future work will focus on further application in the proposed use cases and the extension of the presented method to non-convex OPF formulations.

### v2

@@ V Numerical Case Study
All numerical results in this section are performed on a system with an eight core and 16 thread Ryzen 9800X3D CPU paired with 64GB of 6000MHz RAM. Provided computation times are the average of ten executions. We use JuMP.jl [undefac] and MOSEK [undefad] to solve optimization problems. For convex sets, Polyhedra.jl allows to extract the set directly form the JuMP.jl model as polytope in H-representation. For the problem formulation, we consider LinDistFlow with loss linearization as elaborated in Section II. This allows to achieve a convex approximation while maintaining the crucial information of reactive power and voltage magnitude in the DSO grid model. As the underlying problem, we consider a multi-stage Optimal Power Flow (MS-OPF) problem, which allows flexibililty aggregation across multiple time steps and serves to evaluate computational scalability for horizons of up to $N=96$.
@@ V-A Comparison of Computation Times
First, we consider a small-scale, 4-bus radial distribution grid adapted from MATPOWER, called case4dist [undefae]. The grid has been extended to include two additional renewable generators at buses three and four, thereby increasing the system’s overall active and reactive power flexibility. This small test case allows the comparison of the presented method with the baseline approach of polytope projection, which can still be evaluated over a short varying time horizon of the MS-OPF. While this projection-based approach does not scale well to larger grids, it remains suitable for evaluating performance in this compact setting.
TABLE I: Computation times of the presented method and polytope projection on the 4-bus distribution grid model for up to $N=4$ timesteps. Fastest computation times in bold.
@@
V-B Evaluation on 15-bus Feeder
Next, we analyze how the presented method performs on a medium-sized distribution grid model.
@@ V-B Evaluation on 15-bus Feeder
Figure 4: Scaling of dimensionality and number of constraints of the 15-bus grid model with time horizon length NN.
[…]
@@
First, we demonstrate the equivalence of the reformulation of the feasible set as $\mathcal{CZ}$. Figure 5 shows that approximating the variable bounds does not impact the accuracy of the resulting $\mathcal{CZ}$. The FOR modeled as $\mathcal{CZ}$ $\mathcal{F}_{\mathcal{CZ}}^{LDF-LL}$ is equivalent to the respective FOR modeled as polytope $\mathcal{F}^{LDF-LL}$. As described in Subsection III-B, this holds as long as the bounding zonotope over-approximates the polytope and all constraints defining the polytope are added to the $\mathcal{CZ}$ representation in the next step of the conversion. Hence, all following computations use sufficiently large variable bounds to decrease computation time. For a comparison of the approximation quality of $\mathcal{F}^{LDF-LL}$ with the true flexibility region we refer to [undefr].
Table II shows the respective timeseries of required computation times for the reformulation of the feasible set as $\mathcal{CZ}$. For the given results, the constraint intersection is performed in parallel. The results clearly show that despite the fact that the computation time scales noticeably, it remains suitable for offline computation on the used test problem, even for a considered time horizon length of $N=96$. Further, $\mathcal{CZ}$ projection can be performed in sub-second times on the 15-bus distribution grid case, even for large time horizons $N$. This highlight the suitability for the presented method to reduce complexity in grid operation via aggregation. Paired with acceptable conversion computation times, this allows the application of aggregation to considerably larger problem sizes, by using the presented method to model the feasible sets of the sub problems as $\mathcal{CZ}$.
@@ VI Conclusion and Future Work
This paper addresses the central computational bottleneck arising in projection-based determination of the FOR for affine grid models via tools of computational geometry. By utilizing $\mathcal{CZ}\mathrm{s}$ as beneficial set representations we can significantly reduce the computation times required for flexibility aggregation via set projection of convex feasible sets while achieving the same accuracy of original polytope based approaches. Hence, this approach combines the computational efficiency of zonotope based approximations for set projection with the capability to model non-symmetric polytopes. The computational efficiency of the presented method allows to perform time-dependent flexibility aggregation in the setting of multi-stage OPF and on large-scale problems without loss of accuracy. Further we show that most of the computation can be performed offline and set modifications can be considered efficiently.
Future work will focus on further application in the proposed use cases and a more detailed numerical comparison of the the representational accuracy to alternative approaches like original zonotope projection. Additionally future work will focus on the extension of the presented method to non-convex OPF formulations as well as the charging-discharging efficiency of batteries and their nonlinear dynamics.